OpenStudy (anonymous):

Which ratio forms a proportion with 15/70? A. 45/210 B. 30/36 C. 25/80 D. 3/15 im confused, the answer I came up with is 15/70 = 3/14? am I doing something wrong, or is there just a typo?

OpenStudy (anonymous):

15/70 = 3/14, but it is also equal to one of the other answer choices.

OpenStudy (anonymous):

im thinking A?

OpenStudy (anonymous):

Yep. 15*3=45, 70*3=210

OpenStudy (anonymous):

yes its A if you multiply it by 3 you will get the same ratio.

OpenStudy (anonymous):

okay! thanks so much! :)